Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

die probability

Expert replies
by matt1234 » Tue Dec 13, 2016 4:23 pm
you roll a die 5 times, what is the probability that you get a 6 exactly once


Please help thank you!!!!
Join the discussion
Source: — Problem Solving |

by Jay@ManhattanReview » Wed Dec 14, 2016 4:18 am
matt1234 wrote:you roll a die 5 times, what is the probability that you get a 6 exactly once


Please help thank you!!!!
The probability of getting a '6' = 1/6;

=> The probability of NOT getting a '6' = 5/6; (we can get any number other than the '6': 1, 2, 3, 4, or 5

=> (Probability of getting a '6' on the first throw) * (Probability of NOT getting a '6' on the subsequent four rolls) = (1/6)*(5/6)*(5/6)*(5/6)*(5/6)

= (1/6)*(5/6)^4

= 5^4/6^5

The '6' not necessarily should show up on the first roll, it can show up on 2nd, 3rd, 4th, or 5th roll too.

=> Probability that you get a 6 exactly once = 5C1 * (5^4/6^5) = 5 * 5^4/6^5 = (5/6)^5

Hope this helps!

-Jay

_________________
Manhattan Review GMAT Prep

Locations: New York | Singapore | Doha | Lausanne | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

by [email protected] » Wed Dec 14, 2016 10:11 am
Hi matt1234,

In these types of prompts, if you don't immediately 'see' the elegant approach to getting to the correct answer, it often helps to 'map' out what you're looking to calculate. In that way, you might be able to discover a short-cut that can save you some time as you're working through the math.

We're asked to roll a die 5 times and we're asked for the probability of getting EXACTLY ONE 6. Let's start by mapping out one possible option:

First roll is 6, the last four rolls are 'not 6': (1/6)(5/6)(5/6)(5/6)(5/6) = (5^4)/(6^5)

Of course, we could end up getting the lone 6 on the SECOND roll, with the other 4 rolls being 'not 6'; that would look like this....

(5/6)(1/6)(5/6)(5/6)(5/6) = (5^4)/(6^5)

Notice how the outcome is exactly the SAME. This proves that we don't have to redo any additional calculation. There are 5 'spots', so there are 5 options to end up with just one 6. Thus, we just have to multiply the above calculation by 5....

(5)(5^4)/(6^5)

The numerator can then be 'combined'....

(5^5)/(6^5)

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Thu Dec 15, 2016 4:50 pm
matt1234 wrote:you roll a die 5 times, what is the probability that you get a 6 exactly once
The probability of rolling a 6 is 1/6, and the probability of NOT rolling a 6 is 1 - 1/6 = 5/6.

So, if we are rolling a 6 exactly once in 5 tries, we are rolling a 6 in one of the 5 tries and a number other than 6 in the other 4 tries. We can calculate this as follows (note: S = rolling a 6, NS = not rolling a 6):

P(S x NS x NS x NS x NS) = 1/6 x 5/6 x 5/6 x 5/6 x 5/6 = 625/7776

However, there are 5 ways in which we could roll exactly one 6:

S-NS-NS-NS-NS

NS-S-NS-NS-NS

NS-NS-S-NS-NS

NS-NS-NS-S-NS

NS-NS-NS-NS-S

Thus, the final probability of rolling just one 6 is 625/7776 x 5 = 3,125/7,776

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by crackverbal » Thu Dec 15, 2016 8:56 pm
Hi Matt,

Pretty much everyone has posted what could be different ways of approaching this question.

I will give my 2 cents to this question.

These types of question can be solved using combinatorial approach, but the below rule would help you to approach it better.

Probability of Complex event = Product of individual events * arrangements

Here its rolling the die five times and getting six exactly once.

So according to the above rule it is, (six)(Not six) (Not six) (Not six) (Not six) * arrangement

P(getting six) = 1/6

P(not getting six) = 5/6

So, 1/6 * 5/6 * 5/6 * 5/6 * 5/6 * arrangement.

Arrangement can be done many ways, already few have posted different ways (like selections, listing..) of arranging.

Another alternative way is,

Consider this as a word,

SNNNN,

S – Six , N – Not Six

So SNNNN can be arranged in 5!/4! Ways. That is 5 different ways.

If you are not aware of the above arrangement,

Let’s say for an example, Given a word AISHA – different ways of arranging it is 5!/ 2!. Here the numerator is
arranging of 5 letters and denominator is dividing the repetitions.

So the final answer would be 5^ 5/ 6^5.

So if you look at it finding the probability is similar but arrangements can be done in multiple ways.

So you can chose best way you are comfortable with.

Hope this is clear.
Join Free 4 part MBA Through GMAT Video Training Series here -
https://gmat.crackverbal.com/mba-throug ... video-2018

Enroll for our GMAT Trial Course here -
https://gmatonline.crackverbal.com/

For more info on GMAT and MBA, follow us on @AskCrackVerbal
Join the discussion